A typical incandescent bulb has an efficiency of about 10%. What current is needed for a bulb with a resistance of 1.5×102Ω to output power of 6.0 W?


1
Expert's answer
2022-03-14T13:06:20-0400

Solution;

"P=I^2R"

"6=I^2\u00d71.5\u00d710^2"

"I^2=0.4"

"I=0.6324A"